OpenStudy (anonymous):

If p^6=r and p^9=s, which of the following must be equal to p^12 ? A)2s-r B)2s/r C)s+s/r D)s^2/r E)s^9/r

OpenStudy (anonymous):

how did you do it

OpenStudy (anonymous):

becuase i tried and i couldnt get anything right

OpenStudy (anonymous):

can you explain what you did but i dont know what you did

hero (hero):

Yeah, in vyew

OpenStudy (anonymous):

vyew??

OpenStudy (anonymous):

I think hes saying "View"

OpenStudy (anonymous):

Hero i understand it now

OpenStudy (anonymous):

even though you didnt finish explaining it

OpenStudy (anonymous):

I THINK hes working on it so just relax..

OpenStudy (anonymous):

@Hero i think u r wrong.... 2X =X+X NOT X*X......

hero (hero):

s = p^9 s + s/r = p^9 + p^9/p^6 = p^9 + p^3 = p^3^3 + p^3^1 = p^3(3+1) = p^3^4 = p^12 Therefore, C

OpenStudy (anonymous):

\[a^{n}a^{m}= a^{n+m}\] \[\frac{a^{m}}{a^{n}}=a^{m-n}\] use these properties to solve this question!!

hero (hero):

If you have the solution, post it

OpenStudy (anonymous):

\[p^{9} p^{9}=p^{9+9}=p^{18}=s^{2}\] \[\frac{s^{2}}{r}=\frac{p^{18}}{p^{6}}\] \[\frac{s^{2}}{r}=p^{12}\]